Fermi National Laboratory
 

Operation Summary for June 2, 2010 - June 4, 2010

 

Wednesday June 2

The day shift began with operators preparing the Tevatron (TeV) for access, with Operations monitoring the Recycler stash (422.9E10), with the Antiproton Source (Pbar) stacking (24.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

During the day shift, the CDF, TeV A-E sectors, and Transfer Hall remained in access for necessary repairs. The TeV remained in quench recovery until noon and then operators began the TeV turn-on and orbit smoothing. The A1 cryo system cooldown progressed slowly. RF personnel replaced a power amplifier for a TeV RF station (TRF4).

During the evening shift, Operations put the TeV into shot setup at 4:23 PM and established store 7867 at 5:24 PM with an average initial luminosity (AIL) of 357.33E30. The store aborted at 9:18 PM due to a TeV B11 vertical separator spark; it was a clean abort with no quench. Operations put the TeV into shot setup at 10 PM and established store 7868 at 11:06 PM with an AIL of 293.33E30. Store 7868 was a 36x32 due to the MI Power Supply interlock loop dropping out and with the MI-30 QD VCB opening up moments prior to the first antiprotons transfer. (36x32: The first number, 36, is for bunches of protons and the second number, 32, is for bunches of antiprotons.)

Thursday June 3

The midnight shift began with Operations monitoring store 7868 (215.92E30) and the stash (67.0E10), with Pbar stacking (8.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported some very minor downtimes due to the following: a NuMI beam permit trip caused by a MI incomplete RF gap envelope (no stations tripped), due to a an ion pump (IPRF14) fault that tripped off two Booster RF stations (BRF1 and BRF2), due to a Linac RF station (LRF1) power amplifier crowbar, and finally due to a Klystron RF station (KRF5) gun spark.

During the day shift, Operations terminated store 7868 at 12:03 PM, put the TeV into shot setup, and established store 7869 at 1:34 PM with an AIL of 330.9E30. Operators halted the Proton Source beam for a load-shed test at 1:30 PM. Operations turned off Booster, MI, Debuncher, and several beamlines to shed another 15MW of power for the load-shed test; the test lasted a couple of hours. RF personnel accessed Booster for RF work.

During the evening shift, Operations recovered from load-shed test. The RF technicians came out of Booster at 5:15 PM. Operations had reestablished beam to all users by 5:49 PM.

Friday June 4

The midnight shift began with Operations monitoring store 7869 (84.38E30) and the stash (271.2E10), with Pbar stacking (26.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7869 at 7:47 AM. The SDA/SAR database computer needed a reboot. Operations held off shot setup until controls personnel completed the reboot.

The Plans for Friday and the Weekend

The plans for today and for the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/04/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory